Anhidrotic ectodermal dysplasia is a rare X-linked recessive disorder that results in the absence of teeth, hair, and sweat glands. If the only recent case of this disease in a womans family is an affected brother, what is the probability that her two year old boy will also display symptoms of anhidrotic ectodermal dysplasia?
Answer #1

We will represent X = normal chromosome

Xd= X chromosome having allele for dysplasia

Genotype of woman's brother Xd Y as he is affected. We know that boy gets his X chromosome from mother and Y from father. This means that mother carried Xd allele but as she is normal second X chromosome is normal. Woman's mother is carrier and her father is normal (as he do not show the trait).There is 1/2 possiblity that the woman is carrier. Further if she is carrier there is 1/2 possibility that she will pass the mutated X chromosome to her boy.

Probability of her boy will be affected= 1/2*1/2 = 1/4